I have absolutly NO idea how to do proofs. We just started them today so they are basic 2 section proofs. Are there any basics I should learn before I attempt these?: I have absolutly NO idea how to do proofs. We just started them today so they are basic 2 section proofs. Are there any basics I should learn before I attempt these?
Answer by checkley77(3848)

X^2+8X+16=0
(X+4)(X+4)=0
X+4=0
X=-4 ANSWER.
PROOF:
-4^2+8*-4+16=0
16-32+16=0
32-32=0
0=0
-------------------------------------------------
X^2-5X+6=0
(X-2)(X-3)=0
X-2=0
X=2 ANSWER.
X-3=0
X=3 ANSWER.
PROOFS:
2^2-5*2+6=0
4-10+6=0
10-10=0
0=0
3^2-5*3+6=0
9-15+6=0
15-15=0
0=0